LHC 2012 - latest news

Week 10 Checkout coordinators: R. Giachino, M. Albert, J. Wenninger




Link to report from morning meeting

Monday 12th Mar

  • Access.

8:30 meeting

Sunday 11th Mar

  • Morning: access for ATLAS cooling until ~15:00.
  • Afternoon and evening : TOTEM and ALFA interlock tests, including SMP parameters.
  • Commissioning of RQT13.L8B1 to 440 A (was 400 A).
  • Squeeze test to 0.6 m - OK !
  • Squeeze attempt for high beta (90 m) failed due to current limit in RQF (6400 A, 6415 A required).

9:00 meeting

Saturday 10th Mar

  • LBDS tests.
  • 10:30 Beam to TI8 downstream TED.
  • Problem with injection timing tables - fixed by expert. Some duplicate hidden tables in the MTG HW.
  • 14:00 LBDS TSU tests finished.
  • 15:00 Beam down to TI2 TED.
  • MKE4 waveform scan in TI8 with probe bunches.
  • MKI8 pulsing and tested.
  • 23:00: MKI and LBDS tests finished and almost completed. Delay of ~160 ns on the MKIs to to adjusted.
  • Overnight - run again through nominal sequence. Issue with RQT13.L8B1 were the IPNO is set to 400 A while 405 A are required for the squeeze.

9:00 meeting

Friday 9th Mar

  • TOTEM interlock tests.
  • 20:00 All valves open, LHC closed.
  • 21:30 : beam dumped armed for the first time in 2012 !
  • Ramp with LBDS armed.
  • 22:00 Machine at 4 TeV with armed beam dump.
  • BETS interlock tests at 4 TeV on RBs, Q4s, MSDs.
  • Cycle tests - so issue with settings incorporation.

8:30 meeting

Thursday 8th Mar

  • LBDS BETS tests. Machine cycled with most PCs, including the squeeze.

8:30 meeting

Wednesday 7th Mar

  • LBDS BETS test on RBs, Q4s and MSD in 6. OK.
  • Beta* reconstruction and distribution test (SIS+SMP+MTG). Ok from injection to end of squeeze (0.6/3/0.6/3m).
  • S23 through ramp and squeeze (except QTD/F) OK.

8:30 meeting

Tuesday 6th Mar

  • LBDS BETS test on RBs, Q4s and MSD in 6. RB45 missing. Calibration curves for Q4 and MSD were found clamped.

8:30 meeting

Monday 5th Mar

  • Checkout tests

8:30 meeting